OpenStudy (anonymous):

IVE BEEN ON THIS FOR 2 WEEKS, I NEED HELP! PLEASE EXPLAIN! MEDAL AS REWARD! Match each one with its rise over run. 1. (2, -3), (-1, 4) 2. (-2, 3), (-1, 4) 3. (2, -3), (1,4) 4. (-2, 3), (1, 4) ------- Answer choices: 1 -7/3 -7 1/3

OpenStudy (btaylor):

Do you know what rise and run are? rise = change in y = y2-y1 run = change in x = x2-x1 Call one coordinate 1, and the other 2. Label each value (x1, y1) and (x2, y2). Then, find the rise and run. rise/run = slope = your answer.

OpenStudy (anonymous):

@BTaylor I did that but then I got something so off and now I am frustrated and dont know what to do anymore.

OpenStudy (anonymous):

graph the points on a graph and then count the times it rises up and how many times it runs left or right from your first set of points to the next set. then that will be rise/run

OpenStudy (anonymous):

@sehun after that do I add, or divided the rise and run?

OpenStudy (campbell_st):

an easy way to do it is do a simple plot of the points